OpenStudy (mathlegend):

Barry is four years older than his brother Cole, who is four years older than their sister Darcy. If the sum of their three ages is 60, how old is Barry?

OpenStudy (mathlegend):

C + 4 = B D + 4 = C D + 8 = B B + C + D = 60

OpenStudy (mathlegend):

I tried using substitution but that didn't work...

OpenStudy (bibby):

can't you just replace them all in terms of D? D+8+D+4+D=60 3D+12=60 3D=48 D=16

OpenStudy (mathlegend):

Why did you replace all the terms with D? That did work though.

OpenStudy (bibby):

I can't pinpoint why. I was doing things in my head and that worked out. Apologies for how vague that is. I guess if substituting in for one variable doesn't work, try the others.

OpenStudy (mathlegend):

The question is actually multiple choice but I didn't know what to do without plugging in the answers. Thanks.

OpenStudy (mathlegend):

Okay, I think I have an explanation of why this works...

OpenStudy (mathlegend):

We chose to solve for Darcy the youngest... the variable D was in two out of the three equations... D + (D+4) + (D+8) = 60 What each term Represents: D = Darcy D+4 = Cole D+8= Barry

OpenStudy (bibby):

lemme try doing it for the oldest then C + 4 = B D + 4 = C D + 8 = B B + C + D = 60 B + (B+12) + (B+8) = 60 3b +20 = 60 3b = 40 bleh

OpenStudy (mathlegend):

No, the oldest Barry would work out like this...

OpenStudy (mathlegend):

B - 4 = C B - 8 = D C - 4 = D We have Barry's variable in two of the equations... so lets override the rest like you said... B + (B-4) + (B- 8) = 60 3B - 12 = 60 3B = 72 B = 24

OpenStudy (mathlegend):

Thanks again @bibby

OpenStudy (bibby):

no problem haha